OpenStudy (agl202):

Help... The formula for the circumference of a circle is C = 2πr, where r is the radius. Rearrange the formula to solve for r and select the correct option below. 1.) r = C over 2 pi 2.) r = 2πC 3.) r = C − 2π 4.) r = 2 pi over C

rishavraj (rishavraj):

wht do u think???

rishavraj (rishavraj):

its same like a = b/x b = ax

OpenStudy (agl202):

I think 4.

rishavraj (rishavraj):

why so??? o_O

rishavraj (rishavraj):

if r = 2 pi C then it means C = r/(2 pi)

OpenStudy (triciaal):



OpenStudy (triciaal):

r is on the right what do you have to do to leave r by itself?

OpenStudy (triciaal):

this is an equation. whatever you do to one side also do to the other to maintain the balance

OpenStudy (agl202):

no

rishavraj (rishavraj):

u remember 2x = 8 x = 8/2 x = 4

OpenStudy (agl202):

yes

rishavraj (rishavraj):

so then i guess u can do this one :))

OpenStudy (cggurumanjunath):

option 1

OpenStudy (spring98):

wow